OpenStudy (anonymous):

lim x->0 sin3x sin5x/x^2

OpenStudy (owlfred):

\[\sin (3x)\sin(5x) = \frac{\cos(2x) - \cos(8x)}{2}\]

OpenStudy (owlfred):

Or just use L'Hospital's.

OpenStudy (anonymous):

ah oh okay got it thank you

OpenStudy (owlfred):

:)

OpenStudy (xapproachesinfinity):

you can also separate that x^2=x*x then multiply and divide by 3 as well as 5 then you have the usual limit sinx/x at zero which is 1 and the numbers that you multiplied by are what remains